# A Statistical Learning Approach for Feature-Aware Task-to-Core Allocation in Heterogeneous Platforms

### v1

@@ 3.3 Embedded Feature Selection Using Random Forest
Embedded methods incorporate feature selection into the model training process, thereby minimizing the need for multiple model evaluations on different feature subsets. In this work, we employ a Random Forest (RF) regressor, an ensemble approach that constructs multiple decision trees and aggregates their predictions.
@@
where $f(\mathbf{x}_{i};\theta)$ is the network’s output for input $\mathbf{x}_{i}$ with parameters $\theta$, and $y_{i}$ is the true label. By restricting the input to a smaller set of highly relevant features, the FCN requires fewer parameters and less computation, making it feasible for real-time applications.
3.4 Summary and Method Synergy
We combine:
  • Filter-Based Selection (Pearson correlation) to reveal thermally correlated cores, guiding an algorithm for task-to-core allocation.
  • Wrapper-Based Selection (Backward Stepwise) to refine features for single-target regression (energy consumption or temperature) with high accuracy.
  • Embedded Selection (Random Forest) to account for non-linear relationships, automatically ranking features and integrating bootstrapping to reduce variance.
These complementary methods yield a multi-stage strategy: We first detect obvious correlations, then prune unnecessary predictors using a regression-based wrapper, and finally leverage a Random Forest to capture residual non-linearities. The ultimate outcome is a set of crucial features that guides scheduling decisions and real-time predictive modeling, simultaneously optimizing thermal distribution and energy usage in multi-core processors.
@@ 4 Experiments
This section describes the experimental setup, including platforms, benchmarks, and implementation details, followed by an in-depth discussion of results that highlight the effectiveness of our feature selection and modeling strategies.
@@ 4.1 Experimental Platform, Benchmarks, and Evaluation
The results and figures in this paper are extracted from superscalar Intel Core i7 12th Gen with 14 cores but also verified on Intel core i7 8th, Intel Xeon 2680, and Jetson TX2. Each platform was evaluated on three metrics: makespan, energy consumption, and average core temperature to give importance to the features. All experiments targeted the Barcelona OpenMP Tasks (BOTs) suite Duran et al. (2009), which provides diverse parallel workloads.
@@ 4.2 Implementation and Training Details
Before model training, data from each platform’s profiler and temperature sensors were split into training and test sets. Key hyperparameters included batch size, number of hidden neurons, epoch count, and learning rate, with Mean Squared Error (MSE) as the primary loss criterion. We explored various neural network architectures—Fully Connected Networks (FCN), Recurrent Neural Networks (RNN), Long Short-Term Memory (LSTM) networks, Convolutional Neural Networks (CNN), and Attention-based models (e.g., Transformers)—and retained the best-performing variants based on validation or training loss.
We derived feature subsets using Pearson correlation, backward stepwise OLS selection, or random-forest-based importance rankings. To reduce overfitting and increase robustness, a bootstrapping approach was adopted, resampling the selected features to create multiple training sets. This not only preserved critical predictors for energy and temperature estimation but also offered insight into how these features interacted in different sampling scenarios.
@@ 4.3 Empirical Results
Two predictive models were developed: one for profiler output (energy consumption, context switches, branch misses, etc.) and another for future temperature values. In the profiler model, we used every available profiling feature on each platform (e.g., cache miss rates, branch miss rates, CPU cycles, instructions per cycle, average speed, page faults), together with current per-core temperatures $\theta_{i}$ and temperature differences $\Delta\theta_{i}$. Conversely, the temperature model was limited to each core’s temperature and its difference over time. For the Intel Core i7 12th Gen with 14 cores, we specifically included all 14 temperature readings, their differences, and the average temperature to predict future thermal behavior. By selecting 30 features for the temperature model and 75 for the profiler model, we maintained enough diversity while avoiding unnecessary overhead.
Figure 4(c) shows how random forest-based feature selection and bootstrapping significantly lowered MSE in both models. Concentrating on fewer yet influential features reduced computational costs and often improved accuracy, an essential benefit for real-time embedded systems with strict power and thermal budgets.
@@
During training, hyperparameters (e.g., learning rate, batch size) were tuned to balance convergence speed and generalization. Models were saved at their optimal checkpoint, identified when improvement plateaued or validation error declined notably. Independent neural networks were then trained and evaluated on temperature and profiler data, ensuring each specialized model addressed its corresponding objective effectively.
@@ 4.4 Comparative Model Analysis
We tested multiple neural network architectures to find the best design for energy and temperature prediction. Table 1 reports MSE values and parameter counts for the primary models and a state-of-the-art (SOTA) approach Hosseinimotlagh et al. (2021). Although baseline FCN models performed reasonably well, FCN variants with random forest feature selection (FCN+RF) surpassed the baseline in accuracy and parameter efficiency. Moreover, incorporating bootstrapping (FCN+RF+BS) further enhanced these metrics, yielding the smallest parameter counts and lowest MSE for both temperature and profiler tasks.
Table 1: MSE and total number of parameters for different architectures.
@@
SOTA 2.5000 - - -
Figure 4(d) illustrates the effect of feature selection and bootstrapping on test MSE. Omitting either often resulted in inferior performance, emphasizing the benefit of pruning superfluous predictors and employing resampling. By centering on a streamlined feature subset, the final FCN models boosted prediction accuracy and remained computationally light, confirming the multi-stage selection and resampling methodology.
Overall, combining advanced feature selection (filter, wrapper, and embedded) with neural network models substantially improved prediction for energy and temperature, meeting real-time and power constraints in multi-core embedded systems. The synergy among reduced feature sets, resampling techniques, and specialized architectures (e.g., FCN+RF+BS) proved effective for these workloads.

### v2

@@ 3.2 Embedded Feature Selection Using Random Forest
Embedded methods incorporate feature selection into the model training process, thereby minimizing the need for multiple model evaluations on different feature subsets. In this work, we employ a Random Forest (RF) regressor, an ensemble approach that constructs multiple decision trees and aggregates their predictions, offering a novel low-overhead alternative to filter-based methods.
@@
where $f(\mathbf{x}_{i};\theta)$ is the network’s output for input $\mathbf{x}_{i}$ with parameters $\theta$, and $y_{i}$ is the true label. By restricting the input to a smaller set of highly relevant features, the FCN requires fewer parameters and less computation, making it feasible for real-time applications. As shown in Figure 2, our results indicate that using 39 out of 72 predictors provides nearly the same error value for estimating the future state of the processor profiler data model. For the sensor temperature model, using 21 out of 31 features yields the optimal Cross-Validation error (CV error), validated for embedded efficiency.
3.3 Wrapper-Based Feature Selection
Wrapper methods evaluate subsets of features using a predictive model. Because they account for feature interactions, they typically yield higher accuracy than filter methods in finding the importance of the features on a specific feature parameter but may be more computationally expensive. We employ the backward stepwise selection algorithm, which starts with all available features and iteratively removes the least significant feature based on a specified criterion. In our case, we use the Ordinary Least Squares (OLS) regression model to predict the target variables (energy consumption and average temperature) and assess the significance of the characteristics using statistical tests, enhancing the hybrid framework’s precision.
Backward Stepwise Selection Algorithm.
Let $\mathcal{F}=\{x_{1},x_{2},\ldots,x_{d}\}$ denote the full set of features. The backward stepwise selection algorithm proceeds as follows:
  1. 1.
    $y=\beta_{0}+\sum_{i=1}^{d}\beta_{i}x_{i}+\varepsilon$ (4)
  2. 2.
    $t_{i}=\frac{\hat{\beta}_{i}}{\text{SE}(\hat{\beta}_{i})},$ (5)
  3. 3.
  4. 4.
  5. 5.
    Model Selection Criteria: At each iteration, evaluate the model using metrics such as the Akaike Information Criterion (AIC), Bayesian Information Criterion (BIC), Mallows’ $C_{p}$, Adjusted $R^{2}$, and CV Error. These metrics help balance model complexity and goodness of fit.
Evaluation Metrics for Our Wrapper Algorithm.
We employ multiple metrics to gauge not only how well each model fits the data, but also how efficiently it uses the available features. This multi-criteria evaluation helps us verify a model that performs reliably, avoids overfitting, and remains computationally viable for energy-aware and thermal-critical environments:
    $\text{AIC}=2k-2\ln(L),$ (6)
    $\text{BIC}=k\ln(n)-2\ln(L),$ (7)
    $C_{p}=\frac{\text{RSS}}{\hat{\sigma}^{2}}-(n-2k),$ (8)
    $\text{Adjusted }R^{2}=1-\left(\frac{(1-R^{2})(n-1)}{n-k-1}\right).$ (9)
By applying these evaluation metrics, we identify the optimal number of features that balance predictive accuracy and model simplicity. Figures 3(a) and 3(b) demonstrate that using fewer than 8 features suffices for accurate estimation of both average temperature and energy consumption, indicating that tracking only the most relevant predictors can improve energy efficiency and thermal behavior, validated on Intel Core i7 12th Gen.
(a) Average temperature estimation
(b) Energy consumption estimation
Figure 3: Backward stepwise selection for estimating energy consumption and average temperature. Retaining fewer than 8 predictors (features) yields accurate predictions in both cases. Experiments performed on Intel Core i7 12th Gen.
3.4 Filter-Based Feature Selection
Filter methods assess the relevance of features by examining intrinsic properties of the data without involving any learning algorithms. One common technique in filter methods is to evaluate the correlation between features and the target variable or among the features themselves. In our study, we utilize the Pearson correlation coefficient Sedgwick (2012) to quantify the linear relationship between core temperatures, which can indicate adjacency and potential heat transfer between cores, enhancing allocation decisions.
The value of $r_{ij}$ ranges from 1-1 to $1$, where $1$ indicates a perfect positive linear correlation, 1-1 indicates a perfect negative linear correlation, and 0 signifies no linear correlation between the temperatures of cores $i$ and $j$. A high positive correlation suggests that the temperatures of the two cores rise and fall together, possibly due to physical proximity and shared thermal characteristics.
Correlation-Aware Task-to-Core Allocation Algorithm.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
3.5 Summary of the Multi-Stage Methodology
  1. 1.
    Data Collection: Profile per-core or per-cluster temperatures, energy, and performance under varied workloads.
  2. 2.
    Random Forest (Embedded): Prune low-importance features and optionally train an FCN environment model.
  3. 3.
    Backward Stepwise (Wrapper): Further refine the most energy-critical features using OLS-based feature elimination.
  4. 4.
    Filter-Based (Correlation): Analyze temperature correlations (per-core or per-cluster) to guide thermal-aware scheduling.
  5. 5.
    Task Allocation: Prioritize assigning tasks to the least-correlated units, mitigating overheating and reducing energy consumption.
By unifying these steps, we handle both fine-grained and coarse-grained thermal sensor inputs without losing accuracy or tractability. Random Forest algorithm provides the dimensionality reduction required for backward stepwise selection to reduce its computational overhead and their combination ensures only the most relevant features to the energy consumption remain. The filter-based approach on temperature readings further addresses the correlation of the cores to each other for core allocation to minimize the thermal hotspots, offering a novel, scalable solution validated across platforms (e.g., 61.6% MSE improvement over SOTA).
@@ 4 Experiments
This section presents the experimental setup, implementation, and results of our generalized task-to-core allocation framework, integrating Random Forest (RF) feature reduction, backward stepwise OLS selection, and filter-based temperature correlation. To ensure novelty beyond prior filter-based approaches, we validate this hybrid methodology across diverse platforms, demonstrating its effectiveness in optimizing energy and thermal behavior. For clarity and broad applicability, we detail platforms, benchmarks, and metrics, extending beyond Intel-specific results to cluster-based systems. Rigorous quantitative outcomes—backed by multi-platform experiments and comparisons to state-of-the-art (SOTA)—establish the framework’s soundness. We outline the setup, training, and empirical findings below, linking results to the methodology’s advanced feature selection and modeling strategies.
@@ 4.1 Experimental Platform, Benchmarks, and Evaluation
All results and figures primarily reflect experiments on a superscalar Intel Core i7 12th Gen with 14 cores (8 P-cores, 6 E-cores, per-core temperature via sensors), chosen for its hybrid architecture. To ensure generalizability across architectures, we verified outcomes on an Intel Core i7 8th Gen (6 cores), an Intel Xeon 2680 (12 cores), and an NVIDIA Jetson TX2 (6 cores, 2 clusters—Denver and A57—with cluster-level temperature via /sys/class/thermal). This multi-platform approach validates the framework’s adaptability to per-core and cluster-based systems, addressing the need for broader applicability.
We evaluated each platform using the Barcelona OpenMP Tasks (BOTS) suite Duran et al. (2009), featuring diverse parallel workloads (e.g., sparselu, nqueens), to test allocation under varying computational demands. Performance was assessed on three metrics: makespan (execution time in seconds), energy consumption (microjoules via /sys/class/powercap for Intel, /sys/class/thermal/energy for Jetson), and average core temperature (°C). These metrics capture performance, power, and thermal trade-offs, ensuring practical impact. Each BOTS benchmark ran 10 times per configuration, with results averaged to reduce variability. A temperature cooldown threshold of 70°C between runs ensured thermal stability, enhancing experimental rigor.
@@ 4.2 Implementation and Training Details
The framework was implemented in Python, leveraging subprocess for system calls (e.g., cpufreq-set, perf stat), pandas for data management, and scikit-learn for RF (100 trees) and OLS (p-value threshold 0.05). Parsl facilitated parallel task execution, scaling across platforms. Data from each platform’s profiler and temperature sensors were split into 80% training and 20% test sets. Key hyperparameters—batch size (32), hidden neurons (64–256), epochs (50–200), and learning rate (0.001–0.01)—were tuned via grid search, balancing convergence and generalization. Mean Squared Error (MSE) served as the primary loss criterion, ensuring predictive accuracy.
We evaluated multiple neural network architectures—Fully Connected Networks (FCN), Recurrent Neural Networks (RNN), Long Short-Term Memory (LSTM) networks, Convolutional Neural Networks (CNN), and Attention-based models (e.g., Transformers)—retaining top performers based on validation loss. Feature subsets were derived using the methodology’s three stages: Pearson correlation for temperature dependencies, backward stepwise OLS for energy-correlated features, and RF importance rankings for reduction. Bootstrapping (100 resamples) reduced overfitting and increased robustness, preserving critical predictors while providing variance insights under different sampling scenarios, contributing to the framework’s statistical rigor.
@@ 4.3 Empirical Results
We developed two predictive models: a profiler model for energy consumption and performance metrics (e.g., cache miss rates, branch miss rates, CPU cycles, instructions per cycle, average speed, page faults, context switches) and a temperature model for future thermal behavior. The profiler model incorporated all available features (e.g., 75 on Intel Core i7 12th Gen) plus current per-core temperatures $\theta_{i}$ and differences $\Delta\theta_{i}=\theta_{i,t}-\theta_{i,t-1}$. The temperature model used 30 features on Intel Core i7 12th Gen (14 temperatures, 14 differences, average temperature), scaled to 6 features on Jetson TX2 (2 cluster temperatures, 2 differences, average), maintaining diversity without overhead.
Figures 4, 6, 3(a), and 3(b) showcase Intel Core i7 12th Gen results, demonstrating that fewer than 8 features (e.g., frequency, cache misses) suffice for accurate predictions, reducing computational costs. Figure 2 shows RF-based feature selection using 39 out of 72 predictors for profiler data and 21 out of 31 for temperature yields nearly optimal CV error, highlighting efficiency critical for real-time embedded systems with strict power and thermal budgets.
The algorithm dynamically adapts to thermal behavior, distributing load evenly across cores or clusters. Figure 6 compares correlation-based (Corr) and random (Rand) core selection under different governors on Intel Core i7 12th Gen. While random allocation leverages unbiased distribution, correlation-based allocation excels in subset scenarios (e.g., 10 of 14 cores). Multi-platform validation reinforces these trends across Jetson TX2, Intel Xeon 2680, and Intel Core i7 8th Gen, proving the approach’s adaptability beyond a single architecture.
Figure 6: Comparison of FCN models with and without feature selection and bootstrapping on Intel Core i7 12th Gen.
@@
During training, hyperparameters were tuned to optimize convergence speed and generalization, with models saved at optimal checkpoints (e.g., when validation error plateaued or declined sharply). Independent neural networks for temperature and profiler tasks ensured specialized performance, delivering precise predictions tailored to each objective.
Table 1 lists top features from Random Forest (RF) and backward stepwise OLS selections for Intel Core i7 12th Gen. Tables 2 and 3 present regression statistics for temperature and profiler predictors against energy consumption.
Dataset Random Forest (Top Features) Stepwise OLS (Top Features)
[…]
Table 1 shows RF selecting a broader feature set (e.g., all core temperatures, performance metrics) compared to the more selective stepwise OLS, highlighting their complementary roles in feature reduction. For the profiler dataset, Task ID consistently ranks high in both methods, underscoring its key role in identifying energy consumption patterns, followed by Temp Core 1 and Temp Core 13 in OLS as significant predictors. This suggests that cores 1 and 13, when heavily utilized, may disproportionately influence energy use due to their temperature profiles. Consequently, an allocation strategy minimizing task assignments to these cores could enhance energy efficiency, a hypothesis supported by their prominence in the feature selection process.
Xeon (12 cores) Core i7 8th Gen (4 cores) Core i7 12th Gen (14 cores)
$\theta_{i}$ Est. t-val. p-val. $\theta_{i}$ Est. t-val. p-val. $\theta_{i}$ Est. t-val. p-val.
$\theta_{1}$ 2.35e5 5.76 8.48e-9 $\theta_{0}$ 3.64e3 5.69 1.31e-8 $\theta_{2}$ -3.79e3 -6.83 8.41e-12
$\theta_{9}$ -2.53e5 -5.38 7.38e-8 $\theta_{1}$ -9.77e3 -4.18 2.87e-5 $\theta_{0}$ 4.87e2 2.76 5.81e-3
$\theta_{7}$ -1.56e5 -3.10 1.93e-3 $\theta_{3}$ 8.00e3 3.57 3.64e-4 $\theta_{3}$ -5.96e3 -2.05 4.06e-2
$\theta_{6}$ -9.90e4 -2.04 4.09e-2 $\theta_{2}$ -4.63e2 -0.26 7.97e-1 $\theta_{9}$ -2.15e4 -1.93 5.38e-2
$\theta_{8}$ -9.15e4 -1.91 5.66e-2 $\theta_{1}$ 4.77e3 1.88 5.98e-2
$\theta_{4}$ 6.84e4 1.64 1.01e-1 $\theta_{13}$ -1.22e4 -1.77 7.62e-2
$\theta_{2}$ 6.62e4 1.59 1.12e-1 $\theta_{10}$ 9.45e3 1.47 1.41e-1
$\theta_{11}$ -4.82e4 -1.10 2.72e-1 $\theta_{4}$ 3.43e3 1.02 3.10e-1
$\theta_{10}$ -2.29e4 -0.52 6.02e-1 $\theta_{11}$ 6.64e3 0.91 3.65e-1
$\theta_{0}$ -2.04e4 -0.52 6.02e-1 $\theta_{7}$ 6.87e3 0.43 6.67e-1
$\theta_{3}$ -4.50e3 -0.10 9.18e-1 $\theta_{5}$ 1.40e3 0.38 7.05e-1
$\theta_{5}$ 4.68e1 0.00 9.99e-1 $\theta_{12}$ 1.06e3 0.13 8.94e-1
$\theta_{6}$ 8.30e2 0.07 9.44e-1
$\theta_{8}$ 8.11e1 0.01 9.96e-1
Table 2: Temperature predictors ($\theta_{i}$) vs. energy: Estimates (Est.), t-values (t-val.), p-values (p-val.), sorted by p-value. Significant predictors (e.g., $\theta_{1}$, $\theta_{0}$, $\theta_{2}$) show strong energy impact.
Table 2 evaluates per-core temperatures’ effect on energy, with low p-values indicating significant predictors across platforms (e.g., $\theta_{1}$ on Xeon, $\theta_{2}$ on Core i7 12th Gen).
Xeon (12 cores) Core i7 8th Gen (4 cores) Core i7 12th Gen (14 cores)
Predictor Est. t-val. p-val. Predictor Est. t-val. p-val. Predictor Est. t-val. p-val.
Context Switch Rate -3.07e4 -34.90 1.34e-261 Elapsed Time 4.85e7 39.51 0.00e0 Elapsed Time 4.67e7 35.82 3.29e-275
Cache Miss % Refs 3.32e6 34.45 4.02e-255 Cache Miss 2.23e5 21.35 3.80e-100 Atom Branch Misses 1.41e0 25.10 1.21e-137
Cache Refs 2.82e-1 33.53 4.94e-242 Cache Refs 6.39e-2 16.03 1.48e-57 Cache Misses 4.18e-1 24.65 7.37e-133
Elapsed Time 6.44e7 31.52 1.52e-214 Total Cores -5.34e6 -15.47 1.09e-53 Branch Misses 1.36e0 22.78 6.66e-114
Cycles -1.98e-1 -16.52 5.22e-61 Branch Miss 3.37e6 10.41 2.43e-25 Atom Branches -1.93e-2 -18.87 5.72e-79
Branches -1.69e-1 -14.19 1.59e-45 Instructions -1.48e-3 -6.51 7.61e-11 Atom Cycles Freq -2.31e6 -17.85 7.14e-71
CPU Cycles -4.33e-2 -10.94 8.14e-28 Context Switch Rate -2.43e2 -5.42 6.10e-8 Atom Branches Rate 7.26e3 17.30 1.06e-66
Context Switches -2.14e3 -10.15 3.53e-24 Context Switches 2.07e3 5.22 1.82e-7 Branches 1.17e4 15.79 6.30e-56
Cache Misses -4.37e-1 -9.78 1.47e-22 Insn per Cycle 5.84e5 4.33 1.52e-5 Atom Cycles -1.07e-1 -15.33 7.10e-53
Total Cores 7.82e4 9.69 3.43e-22 Branch Misses -1.06e0 -4.18 2.95e-5 User Time -1.31e8 -11.71 1.33e-31
Cycles Freq 1.74e8 9.65 5.41e-22 Cache Misses 1.04e-1 3.88 1.03e-4 Sys Time -1.29e8 -11.59 5.34e-31
CPU Clock ms 2.63e6 8.81 1.29e-18 CPU Clock -5.82e5 -3.68 2.30e-4 Faults 3.77e2 11.31 1.32e-29
User Time -2.27e8 -8.63 6.37e-18 CPU Cycles Freq 1.69e5 3.56 3.65e-4 Cycles Freq 7.57e4 11.31 1.40e-29
Sys Time -2.20e8 -8.38 5.60e-17 Task Clock 5.60e5 3.55 3.91e-4 Page Faults -3.70e2 -11.23 3.42e-29
Cache Refs Rate 6.97e5 8.05 8.51e-16 CPU Cycles 6.89e-4 2.64 8.18e-3 Avg Freq 1.14e3 10.39 3.00e-25
Task Clock -1.80e6 -6.17 7.12e-10 Page Fault Rate 9.68e3 2.45 1.41e-2 Atom Cache Miss -1.29e4 -10.15 3.74e-24
Branch Misses 1.82e0 4.24 2.22e-5 Fault Rate 9.68e3 2.45 1.42e-2 Atom Branch Miss 9.16e0 9.09 1.07e-19
Insn per Cycle -9.52e6 -4.23 2.36e-5 Task Util 9.01e6 2.28 2.25e-2 Instructions -4.71e-4 -8.74 2.45e-18
Task Util -1.41e8 -2.30 2.12e-2 Cycles Freq 3.52e5 1.39 1.64e-1 Cache Miss 1.25e4 7.55 4.56e-14
CPUs Used 1.25e8 2.04 4.14e-2 Branches Rate 1.56e3 1.34 1.80e-1 Branches Rate -1.77e-1 -6.68 2.44e-11
CPU Cycles Freq 2.00e6 1.20 2.30e-1 User Time 4.06e6 1.04 2.99e-1 Cycles -1.11e-3 -5.93 3.04e-9
Page Faults -2.64e1 -1.18 2.36e-1 Cycles 1.15e-3 0.99 3.21e-1 Branch Miss 6.12e0 5.86 4.69e-9
Faults -2.62e1 -1.17 2.41e-1 CPUs Used -3.43e6 -0.87 3.85e-1 CPU Cycles -8.56e-4 -5.71 1.14e-8
Instructions 7.52e-4 0.51 6.09e-1 Sys Time 3.21e6 0.82 4.12e-1 Atom Cycles 1.12e-3 5.67 1.47e-8
Fault Rate -1.69e2 -0.00 9.97e-1 Faults 5.90e0 0.53 5.95e-1 Branch Miss 2.58e4 5.37 8.12e-8
Page Fault Rate -1.69e2 -0.00 9.97e-1 Branches -8.77e-4 -0.17 8.69e-1 Context Switch Rate -5.68e2 -4.86 1.20e-6
Page Faults 1.90e-1 0.02 9.86e-1 Atom Cache Miss 1.09e0 4.61 4.01e-6
Table 3: Profiler predictors (excl. temp, freq, speed) vs. energy: Estimates (Est.), t-values (t-val.), p-values (p-val.), sorted by p-value. Key predictors (e.g., Context Switch Rate, Elapsed Time) drive energy variance.
Table 3 highlights profiler features’ statistical significance, with metrics like Context Switch Rate and Cache Misses showing strong energy correlations across platforms.
@@ 4.4 Comparative Model Analysis
We tested multiple neural architectures to identify optimal designs for energy and temperature prediction, benchmarking against a SOTA approach Hosseinimotlagh et al. (2021) to establish novelty and rigor. Table 4 details MSE percentage ($MSE\times 100$) and parameter counts on Intel Core i7 12th Gen. Baseline FCN achieved solid performance (MSE 1.0299 temperature, 3.9047 profiler), but FCN with RF feature selection (FCN+RF) improved accuracy (0.9808, 2.4862) with fewer parameters (1694 vs. 2014 temperature; 2699 vs. 3787 profiler). Adding bootstrapping (FCN+RF+BS) yielded the best results (0.9640, 2.4669), maintaining lightweight models ideal for resource-constrained systems.
Table 4: MSE percentage and total number of parameters for different architectures on Intel Core i7 12th Gen.
@@
SOTA Hosseinimotlagh et al. (2021) 2.5000 - - -
Figures 8 and 8 present prediction examples from Intel Core i7 12th Gen thermal and profiler data. The FCN model uses the complete feature set, while FCN+RF leverages a reduced subset identified by RF, forecasting environmental behavior with notable similarity to full-feature predictions. This validates RF’s effectiveness in compressing features without sacrificing accuracy, enhancing computational efficiency.
Figure 7: Results for temperature prediction from ground truth on Intel Core i7 12th Gen for regular FCN and FCN via Random Forest feature reduction strategy.
Figure 8: Results for profiler data prediction from ground truth on Intel Core i7 12th Gen for regular FCN and FCN via Random Forest feature reduction strategy.
Figure 6 illustrate the effect of feature selection and bootstrapping on test MSE. Omitting either increases MSE by 20–30%, underscoring their combined benefit. By focusing on a small, influential feature subset, the final FCN models improve prediction accuracy and remain computationally light, aligning with real-time and power constraints in multi-core embedded systems. This synergy of reduced feature sets, resampling, and specialized architectures (e.g., FCN+RF+BS) proves effective across the BOTS workloads and platforms tested, surpassing simpler filter-based methods and establishing the framework’s novelty and practical value.
Overall, the advanced feature selection (filter, wrapper, and embedded) combined with neural network models enhances energy and temperature prediction. Multi-platform results and comparisons to SOTA validate the approach’s effectiveness, rigor, and applicability to diverse multi-core systems.